What is the minimum [I−] that must be added to a 1.0×10−3M solution of Pb2+ in order to form a precipitate of PbI2(s)? The Ksp of PbI2(s) is 8.5×10−9. Report your answer with two significant figures.
Answer #1

PbI2(s)   <->   Pb2+(aq)   +   2I-(aq)

Ksp = [Pb2+][I-]^2
8.5*10^-9 = (1.0*10^-3)*[I-]^2
[I-]^2 = 8.5*10^-6
[I-] = 2.9*10^-3 M

Answer: 2.9*10^-3 M